OpenStudy (anonymous):

determine whether y is a function of x. y= radical 1-x

OpenStudy (anonymous):

it is a function if you write it as one \[f(x)=\sqrt{1-x}\]

OpenStudy (anonymous):

y is a function of x if there is exactly one y value for any x you might plug in.

OpenStudy (anonymous):

\[f(x) = \sqrt{1-x}\] \[y=f(x)=\sqrt{1-x}\] \[y ^{2}=1-x\] \[x=1-y ^{2}\] plug the x into \[f(x) = \sqrt{1-x}\] \[f(x) =\sqrt{1-(1-y ^{2})}\] \[f(x) = \sqrt{-y ^{2}}\] \[f(x) = (-y ^{2})^{1/2}\] CMIIW satellite